# RadioDiff-Loc: Diffusion Model Enhanced Scattering Congnition for NLoS Localization with Sparse Radio Map Estimation

### v1

@@ Abstract
Accurate localization of non-cooperative signal sources in non-line-of-sight (NLoS) environments remains a critical challenge with a wide range of applications, including autonomous navigation, industrial automation, and emergency response. In such settings, traditional positioning techniques relying on line-of-sight (LoS) or cooperative signaling fail due to severe multipath propagation and unknown transmit power. This paper proposes a novel generative inference framework for NLoS localization based on conditional diffusion models. By leveraging the physical insight that diffracted electromagnetic energy concentrates near building edges, we develop a sampling strategy that collects sparse received signal strength (RSS) measurements at the geometric vertices of obstacles—locations that maximize Fisher information and mutual information with respect to the unknown source. To overcome the lack of known transmission power, we normalize all sampled RSS values relative to the maximum observed intensity, enabling the construction of a power-invariant radio map (RM). A conditional diffusion model is trained to reconstruct the full RM based on environmental layout and sparse RSS observations. Localization is then achieved by identifying the brightest point on the generated RM. Moreover, the proposed framework is compatible with existing RSS-based localization algorithms, enabling a dual-driven paradigm that fuses physical knowledge and data-driven inference for improved accuracy. Extensive theoretical analysis and empirical validation demonstrate that our approach achieves high localization accuracy with significantly reduced sampling cost, offering a scalable and physically grounded solution for non-cooperative NLoS emitter localization.
@@ I Introduction
The growing deployment of Internet of Things (IoT) systems, smart city infrastructure, autonomous robotics, and emergency services has intensified the demand for high-accuracy localization in complex environments [1, 2, 3]. Applications such as indoor navigation [4], automated guided vehicle (AGV) operations in warehouses [5], and firefighter tracking in hazardous buildings rely heavily on precise position estimation [6, 7]. However, a fundamental challenge in these scenarios is the frequent presence of non-cooperative signal sources—transmitters that do not follow standard localization protocols due to their passive nature or hostile intent [7]. Compounding this difficulty is the predominance of non-line-of-sight (NLoS) propagation, where signal paths are blocked by structures like walls, machinery, or debris, resulting in severe multipath interference and signal attenuation [8]. Traditional positioning systems, such as the global positioning system (GPS) or LoS-based methods, perform well in open areas but fail in environments where direct paths are obstructed [9]. In urban canyons, underground tunnels, or smoke-filled buildings, GPS signals are often unavailable or unreliable. Alternatives like vehicle-to-anything (V2X) communications and roadside unit (RSU) cooperation also suffer in NLoS settings [10]. Similarly, in automated warehouses, AGVs must navigate using signals that are frequently blocked or distorted [8]. These limitations highlight the urgent need for localization techniques that are robust to NLoS effects and independent of signal source cooperation.
Despite its critical importance, accurate localization under NLoS conditions remains a fundamentally challenging problem [2]. These difficulties originate primarily from two intertwined sources: the absence of explicit prior knowledge about electromagnetic (EM) wave propagation in complex environments, and the limited capability to perform measurements in regions where the target emitter resides [11]. First, from a theoretical standpoint, NLoS propagation paths are governed by intricate physical interactions, including diffraction, reflection, and scattering across heterogeneous surfaces [12]. These interactions render the propagation path highly environment-dependent and inherently nonlinear, making it infeasible to describe the signal distribution using tractable, closed-form models. Without a known prior distribution of the signal field, conventional localization techniques based on maximum a posteriori (MAP) estimation become ill-posed, as the likelihood function lacks a meaningful Bayesian prior to regularize inference [13]. This is particularly problematic in obstructed environments, where the dominant contributions to the received signal do not correspond to direct paths, but rather to weak and spatially irregular multipath components. Second, the challenge is further exacerbated by constraints on the spatial domain of measurement [14]. In many practical NLoS positioning scenarios, the target radiation source is confined to a region that cannot be directly sampled, either because the area is inaccessible, such as inside collapsed structures or restricted zones, or because the source is non-cooperative and emits passively [8]. If complete spatial sampling were possible, LoS components could potentially be recovered, eliminating the problem altogether. However, NLoS positioning tasks inherently preclude this possibility. Moreover, non-cooperative sources do not emit structured pilot signals at known intervals, rendering traditional metrics such as time-of-arrival (ToA) or time-difference-of-arrival (TDoA) inapplicable. In such cases, spectrum sensing equipment is typically limited to acquiring only coarse-grained observations, such as received signal strength (RSS), which are further confounded by NLoS distortions [15]. Critically, constructing a meaningful estimate of the source location from RSS data alone demands prior knowledge of the propagation field—knowledge that is neither known a priori nor easily measurable. Moreover, the problem of sampling further complicates the NLoS localization process. Since the propagation paths in NLoS environments are not explicitly known, there exists no principled strategy to determine which regions should be sampled or how many samples are necessary to ensure accurate inference. Existing methods often resort to uniform or heuristic-based random sampling, which is inefficient and may fail to capture critical signal variations induced by environmental geometry [16]. This unpredictability in sampling design introduces significant overhead in terms of measurement cost and deployment logistics, hindering the scalability of NLoS localization in real-world settings
To address the theoretical and practical challenges of NLoS positioning, the concept of a radio map (RM) offers a promising direction [17, 18, 19]. An RM represents the spatial distribution of wireless signal characteristics, such as pathloss or received signal strength, across a geographical area, effectively encoding the influence of environmental geometry on electromagnetic propagation [20, 21, 22]. From a probabilistic perspective, each realization of an RM can be interpreted as a sample drawn from the implicit prior distribution that governs signal behavior in a specific environment. In this context, the RM serves as a surrogate for the otherwise intractable Bayesian prior, enabling the localization problem to be cast within a well-defined inference framework [23]. By preconstructing or learning the RM of an environment, one can use sparse measurements from a sensing device to perform posterior inference on the source location, thus enabling accurate NLoS localization even in the absence of direct path observations. However, the utility of RMs in this framework is limited by a crucial constraint: the RM does not offer an explicit parametric form of the prior distribution [21]. Unlike analytical models, which enable closed-form derivations, the RM is a data-driven, high-dimensional structure whose internal dependencies and correlations are difficult to characterize mathematically. This lack of explicit structure hinders the direct application of classical Bayesian inference techniques, which require known priors to compute posterior distributions. Moreover, the spatial heterogeneity of NLoS propagation—driven by building layouts, material properties, and object occlusions—means that each RM is highly specific to its environment. As a result, generalizing across environments requires the ability to learn and represent this distributional prior in a flexible, data-adaptive manner. In this regard, neural networks, particularly those trained under generative paradigms, offer a compelling solution. By learning to model the underlying spatial structure of the RM, generative neural networks can capture the high-dimensional distribution of signal features in a latent space. This representation not only enables sample generation conditioned on sparse inputs but also provides a mechanism to extract and reuse prior knowledge embedded in historical or synthetic RMs. Thus, the RM transitions from a passive map to an active component in probabilistic inference, enabling learning-based posterior sampling for NLoS localization. This perspective opens the door to a fundamentally new class of methods for wireless positioning, in which prior distributions are not analytically specified but learned implicitly through data-driven modeling.
Motivated by the aforementioned NLoS localization challenges and successes of the RM, this paper proposes a novel framework for non-cooperative NLoS localization and RM reconstruction, grounded in a generative diffusion modeling paradigm. Specifically, we develop a simultaneous localization and mapping (SLAM) methodology based on denoising diffusion models, which learn to generate plausible radio signal distributions conditioned on environmental structure and sparse observations. In the proposed framework, the diffusion model is first pre-trained to learn the prior distribution of RSS values across the environment, represented as radio maps, using only building layout and base station (BS) location as prompts. This training phase embeds the spatial propagation characteristics of the environment into the model weights, effectively capturing the statistical properties of the electromagnetic field without relying on manual prior modeling. During inference, the pre-trained diffusion model is fine-tuned or conditioned using a sparse set of RSS measurements acquired by a mobile agent or spectrum sensing device. This enables the model to perform maximum a posteriori (MAP) sampling of the full RM or the emitter’s location by progressively denoising a random initial guess under the learned generative prior. In this way, the system performs efficient Bayesian inference in a learned signal space, adapting to NLoS distortions and measurement uncertainty in a principled manner. To further address the sampling cost bottleneck, which plagues traditional NLoS localization strategies, we propose two physics-inspired measurement strategies derived from knife-edge diffraction theory. Rather than relying on dense or randomly distributed samples, our method strategically places sensing points at the vertices of building geometries, where diffraction-induced field variation is most informative. These sampling schemes reduce the required number of observations to a value that scales only with the geometric complexity of the environment, not the emitter’s position. Remarkably, in low-complexity environments, the required number of samples can be fewer than one per thousand spatial grid points, yielding a localization and RM reconstruction framework that is both sample-efficient and environment-agnostic. The main contributions of this paper are summarized as follows.
@@
    For the first time, we propose a method that localizes NLoS radiation sources using only sparsely sampled RSS measurements from restricted sensing areas. By reconstructing a dense RM from limited observations, our approach enables accurate localization without requiring full-field coverage or source cooperation.
@@
    Inspired by knife-edge diffraction theory, we design two environment-driven sampling strategies, which are surface-based and vertex-based, that drastically reduce measurement cost. These strategies require only a fixed set of samples determined by the environmental geometry, independent of the emitter’s location.
@@
    By reconstructing a dense RM from sparse, high-information RSS measurements using a conditional diffusion model, the proposed framework enables seamless integration with traditional RSS-based localization methods such as trilateration and fingerprinting. This dual-driven design enhances localization accuracy.
@@ II-A Relateds Works
Recent research in NLoS localization has pursued multiple directions to mitigate the challenges posed by multipath propagation and biased measurements, which remain the predominant sources of error in time-based indoor and urban localization systems. One prominent strand of work reformulates NLoS handling as a pattern recognition problem, leveraging deep learning to classify or regress signal attributes directly from raw observations. Early work by Bregar and Mohorčič demonstrated that convolutional neural networks (CNNs) trained on raw ultra-wideband (UWB) channel impulse responses could effectively infer link state and residual range bias, thereby enabling corrected range measurements that, when passed through a classical weighted least-squares solver, achieved over 70% reduction in median localization error. Building on this foundation, Zhao and Wang employed generative adversarial networks to augment class-imbalanced coal mine datasets, improving CNN-based link classification accuracy to 91.2% and realizing measurable localization gains in field deployments. More recent developments have introduced richer input representations—such as channel-state matrices or CSI-derived images—that allow uncertainty estimation and top-K candidate outputs, laying the groundwork for reliable NLoS handling in safety-critical environments.
@@
A second line of research incorporates lightweight learning modules within probabilistic filtering frameworks to improve resilience under mixed LoS/NLoS conditions. Cheng et al. integrated a directional probabilistic data-association gate with an adaptive particle filter to reject spurious ranges and refine tracking performance; similarly, Kang et al. proposed the NNMML pipeline, where hypothesis testing segregates LoS and NLoS links, and a shallow neural network adaptively corrects estimates from parallel Kalman and unscented Kalman filters. Both approaches underline a key insight: combining explicit link-state discrimination with adaptive filter re-weighting consistently outperforms either strategy alone. In contrast to pre-classification-based methods, optimization-driven techniques have emerged that aim to be intrinsically robust to biased measurements. Liu et al. cast the UWB localization problem as a max–min optimization under a correntropy-based loss function, solved using a continuous-time neurodynamic architecture. Notably, this method avoids reliance on link-state labels or statistical bias models and still achieves superior performance in bias-dominated regimes. These developments signal a shift toward robust, bias-invariant estimation objectives as an alternative to pre-screening measurements.
@@
More recently, geometric learning approaches have sought to model multipath explicitly by embedding ray-based physical constraints. In RayLoc, Han et al. formulate localization as an inverse ray tracing problem, optimizing both transmitter coordinates and environmental geometry via a fully differentiable simulator. The integration of Gaussian-kernel convolution mitigates gradient sparsity, enabling convergence to centimeter-level accuracy and outperforming conventional CSI-based methods. This hybrid treatment of geometry and learning exemplifies a growing trend toward physically informed neural inference in wireless positioning. Although these lines of research represent significant progress, they also expose critical limitations. Many learning-based models are tightly coupled to specific input modalities or hardware (e.g., UWB), while optimization frameworks often suffer from high computational cost or require detailed geometric priors. Moreover, both approaches generally assume access to extensive measurement data, which is impractical in scenarios involving non-cooperative sources or inaccessible environments. In contrast, our work takes a fundamentally different approach by recasting the NLoS localization problem as a generative inference task, where the unknown radiation source location and signal distribution are sampled from a learned conditional distribution. By leveraging diffusion-based generative modeling, we learn the spatial prior distribution of electromagnetic signal propagation from environmental geometry and then perform posterior inference based on sparse RSS observations. This framework not only eliminates the need for pilot signals or pre-classified measurements but also offers theoretical guarantees on posterior recoverability under sparse sampling guided by diffraction-informed heuristics. In doing so, our method bridges the gap between data-driven modeling and physics-aware inference, offering a scalable and generalizable solution for high-precision NLoS localization in real-world, measurement-limited scenarios.
Figure 1: Illustration of the RadioDiff-Loc framework.

### v2

@@ Abstract
Accurate localization of non-cooperative signal sources in non-line-of-sight (NLoS) environments is a fundamental yet unsolved challenge, with critical implications for autonomous navigation, emergency response, and smart infrastructure. Traditional localization techniques—relying on line-of-sight (LoS) paths, structured pilots, or known transmit power—fail in environments characterized by severe multipath propagation, restricted sensing regions, and passive emitters. To address this, we propose RadioDiff-Loc, a novel generative localization framework that integrates conditional denoising diffusion models with physics-guided sampling strategies. The model is first trained to learn the statistical prior of the radio map (RM) conditioned on environmental geometry, and then reconstructs the full signal field from a sparse set of received signal strength (RSS) observations. Localization is performed by identifying the intensity peak of the generated RM. To overcome the lack of transmit power knowledge, we introduce a power-invariant normalization scheme, enabling robust inference from uncalibrated RSS. Furthermore, inspired by knife-edge diffraction theory, we design two geometry-aware sampling strategies—surface-based and vertex-based—that place sensors at locations with maximal information gain. These methods drastically reduce measurement costs by aligning sampling complexity with environmental geometry, not emitter position. Extensive experiments demonstrate that RadioDiff-Loc achieves high localization accuracy using over 10× fewer samples than baseline approaches, offering a scalable, interpretable, and physically grounded solution for non-cooperative localization under NLoS conditions.
@@ I Introduction
The rapid growth of autonomous systems, Internet of Things (IoT) deployments, and smart infrastructure in urban and industrial environments has sharply increased the demand for reliable localization under complex and obstructed conditions [1, 2, 3]. Applications such as indoor navigation in large buildings [4], automated guided vehicle (AGV) coordination in logistics warehouses [5], and firefighter tracking in hazardous zones critically depend on accurate position estimation, where traditional methods fail [6]. In these scenarios, the signal source is often non-cooperative, which lacks structured pilot signals or explicit identifiers, and the dominant propagation is non-line-of-sight (NLoS), due to occlusions by walls, machinery, or debris [7]. Under such conditions, mainstream localization techniques—including global positioning systems (GPS) and vehicle-to-everything (V2X) protocols—become ineffective, as they assume either the presence of a direct signal path or prior knowledge of the transmitter’s behavior. For instance, GPS is unreliable in urban canyons, underground tunnels, and smoke-filled buildings, while cooperative solutions based on roadside units (RSUs) or anchor nodes struggle in dynamic or adversarial environments [8, 9, 7]. These limitations expose a significant gap in existing positioning frameworks: the inability to passively localize the unknown sources in geometrically complex and signal-degraded regions. Addressing this challenge is essential not only for safety-critical applications but also for the broader realization of resilient, intelligent, and autonomous communication systems in 6G and beyond.
Despite its critical importance, accurate localization of non-cooperative emitters in NLoS environments remains an open and fundamentally ill-posed problem [2]. The core difficulty lies in two coupled limitations: (i) the absence of tractable priors describing electromagnetic (EM) wave propagation in complex environments [10], and (ii) the inability to perform direct measurements in regions where the target emitter resides [6], which are shown in Fig. 1. First, unlike LoS scenarios where signal behavior can be approximated with simple geometric models, NLoS propagation is governed by complex physical interactions, including diffraction, reflection, and scattering, across heterogeneous surfaces and materials [11]. These interactions cause nonlinear, spatially irregular signal distributions that defy analytical modeling. As a result, classical Bayesian approaches, such as maximum a posteriori (MAP) estimation, become ill-defined due to the lack of a meaningful prior distribution over the signal field [12]. Second, in practical deployments, measurement access is often limited by environmental constraints: the emitter may be located inside a collapsed building, a restricted area, or a dynamically changing zone, rendering direct sampling infeasible [7]. This is further compounded by the non-cooperative nature of the source, which emits passively and lacks synchronization or structured pilots. In such cases, the only available information is coarse-grained received signal strength (RSS) measurements [13], which are severely distorted by multipath propagation and sensitive to both geometry and material properties. Moreover, without knowledge of the transmitter’s power, even the scale of RSS values is ambiguous, further complicating inference. Finally, the lack of explicit propagation paths means that no principled sampling strategy exists to determine where and how to collect measurements for reliable localization. Existing methods often resort to uniform or heuristic-based sampling, which fails to capture the critical variations induced by environmental geometry, resulting in poor accuracy and high deployment cost. These challenges collectively underscore the need for a fundamentally new paradigm that can handle uncertain physics, sparse observations, and adversarial conditions in a unified and scalable manner.
Figure 1: The illustration of the NLoS localization based on RSS information in the allowed area. (a) shows the full RM where the radiation source can be easily obtained from the RM; (b) is the partial RM where the RSS information in the restricted area is unknown; (c) is the random sampling result in the allowed area; and (d) is the diffraction inspired sampling method.
To overcome the lack of analytical models for signal propagation in NLoS scenarios, the concept of a radio map (RM) has emerged as a powerful alternative for enabling data-driven inference [14, 15, 16, 17, 18]. An RM captures the spatial distribution of wireless signal characteristics—such as pathloss or received signal strength—across a geographical region, effectively embedding the influence of environmental geometry, material properties, and electromagnetic interactions into a dense signal field [19]. From a probabilistic standpoint, each RM can be interpreted as a sample from an implicit prior distribution that governs how electromagnetic energy propagates in a specific environment [12, 18]. This perspective opens the door to recasting the localization problem as a posterior inference task, where sparse RSS measurements serve as observations, and the RM acts as a prior for the otherwise intractable Bayesian prior. If this prior distribution could be explicitly modeled or sampled from, it would enable principled estimation of the emitter’s position, even in the absence of LoS paths. However, RMs are inherently high-dimensional, nonparametric structures, whose internal dependencies are shaped by complex environmental factors [20]. As such, they lack closed-form parametric representations, making them incompatible with conventional inference techniques that require tractable priors. Furthermore, because each RM is specific to its underlying environment, it is essential to develop methods that can learn and generalize the prior distribution over RMs, conditioned on observable environmental features such as building layouts or base station positions [17]. This motivates the use of generative models, which are capable of learning high-dimensional data distributions and capturing the latent structure of signal fields in a flexible, environment-adaptive manner [21]. In this way, the RM transitions from a static visualization tool to an active component of the inference pipeline—serving as the learned, physically grounded prior that enables accurate NLoS localization from sparse and noisy measurements.
While learning-based methods can reconstruct the radio map from sparse observations, their performance is fundamentally constrained by the quality and informativeness of the sampled measurements [18]. In NLoS environments, where signal behavior is highly sensitive to geometry, the RSS samples taken often matter more than how many are taken. To address this, we draw inspiration from knife-edge diffraction theory, which reveals that electromagnetic energy tends to concentrate and vary sharply near geometric discontinuities—particularly building edges and vertices [11]. These regions exhibit high sensitivity to emitter position and orientation, making them ideal candidates for sampling under information-theoretic criteria such as Fisher information and mutual information with respect to the unknown source. Based on this physical insight, we propose two geometry-aware sampling strategies: a surface-based scheme, which samples along structural boundaries, and a vertex-based scheme, which targets architectural corners where field variation is most pronounced. Unlike heuristic or uniformly random sampling, which can miss critical spatial features, our approach places sensors at locations where RSS values are not only more informative, but also more stable under power normalization, enabling inference without knowledge of absolute transmission power. Remarkably, the number of required samples in our approach scales not with the size of the environment or the number of grid points, but rather with its geometric complexity—e.g., the number of vertices—yielding substantial gains in sampling efficiency. In low-complexity settings, accurate localization and RM reconstruction can be achieved with fewer than 1% of all spatial positions measured. This physics-informed sampling paradigm transforms sparse measurement collection from an arbitrary design choice into a principled component of the inference framework, enabling practical deployment in real-world NLoS settings with strict sensing constraints.
Motivated by these challenges and opportunities, this paper proposes RadioDiff-Loc, a novel generative localization framework that integrates conditional diffusion models with physics-guided sampling to enable accurate, sampling-efficient localization of non-cooperative NLoS emitters. Unlike conventional approaches that rely on fully sampled observations or handcrafted priors, RadioDiff-Loc learns the statistical structure of radio maps directly from environmental layouts and sparse RSS inputs, enabling end-to-end inference under severe sensing constraints. To the best of our knowledge, this is the first work to realize dense radio map reconstruction and precise emitter localization using only a limited number of uncalibrated RSS measurements from restricted sensing regions, without access to transmitter power or LoS paths. The main contributions of this paper are summarized as folows.
@@
    We propose the RadioDiff-Loc framework to leverage conditional denoising diffusion models for non-cooperative NLoS emitter localization. It reconstructs dense radio maps and estimates emitter positions using only sparsely sampled RSS data from restricted regions—without requiring LoS paths, structured pilot signals, or known transmit power.
@@
    To address the challenge of unknown transmission power, we introduce a normalization scheme that enables the diffusion model to generate relative RSS distributions. This allows robust inference and localization from uncalibrated measurements, making the system agnostic to source power.
@@
    Inspired by knife-edge diffraction theory, we design two geometry-aware sampling strategies—surface-based and vertex-based—that place sensors at locations with maximal information gain. These strategies reduce the required sampling budget by orders of magnitude and scale with environmental complexity rather than emitter location.
  4. 4.
    By generating dense, physically plausible radio maps, RadioDiff-Loc enables seamless integration with traditional RSS-based techniques such as fingerprinting and trilateration, bridging generative modeling with legacy localization pipelines for enhanced robustness and accuracy.
@@ II-A Relateds Works
Recent research in NLoS localization has pursued multiple directions to mitigate the challenges posed by multipath propagation and biased measurements, which remain the predominant sources of error in time-based indoor and urban localization systems [22, 23]. One prominent strand of work reformulates NLoS handling as a pattern recognition problem, leveraging deep learning to classify or regress signal attributes directly from raw observations [24, 25, 26]. Early work by Bregar and Mohorčič demonstrated that convolutional neural networks (CNNs) trained on raw ultra-wideband (UWB) channel impulse responses could effectively infer link state and residual range bias [24], thereby enabling corrected range measurements that, when passed through a classical weighted least-squares solver, achieved over 70% reduction in median localization error. Building on this foundation, Zhao and Wang employed generative adversarial networks to augment class-imbalanced coal mine datasets [26], improving CNN-based link classification accuracy to 91.2% and realizing measurable localization gains in field deployments. More recent developments have introduced richer input representations—such as channel-state matrices or CSI-derived images—that allow uncertainty estimation and top-K candidate outputs, laying the groundwork for reliable NLoS handling in safety-critical environments.
@@
A second line of research incorporates lightweight learning modules within probabilistic filtering frameworks to improve resilience under mixed LoS/NLoS conditions [27, 28, 29]. Cheng et al. integrated a directional probabilistic data-association gate with an adaptive particle filter to reject spurious ranges and refine tracking performance [27]; similarly, Kang et al. proposed a novel pipeline, where hypothesis testing segregates LoS and NLoS links, and a shallow neural network adaptively corrects estimates from parallel Kalman and unscented Kalman filters [30]. Both approaches underline a key insight: combining explicit link-state discrimination with adaptive filter re-weighting consistently outperforms either strategy alone. In contrast to pre-classification-based methods, optimization-driven techniques have emerged that aim to be intrinsically robust to biased measurements [31, 32]. Liu et al. cast the UWB localization problem as a max–min optimization under a correntropy-based loss function, solved using a continuous-time neurodynamic architecture [32]. Notably, this method avoids reliance on link-state labels or statistical bias models and still achieves superior performance in bias-dominated regimes. These developments signal a shift toward robust, bias-invariant estimation objectives as an alternative to pre-screening measurements.
@@
More recently, geometric learning approaches have sought to model multipath explicitly by embedding ray-based physical constraints [33, 34]. In RayLoc [34], Han et al. formulate localization as an inverse ray tracing problem, optimizing both transmitter coordinates and environmental geometry via a fully differentiable simulator. The integration of Gaussian-kernel convolution mitigates gradient sparsity, enabling convergence to centimeter-level accuracy and outperforming conventional CSI-based methods. This hybrid treatment of geometry and learning exemplifies a growing trend toward physically informed neural inference in wireless positioning. Although these lines of research represent significant progress, they also expose critical limitations. Many learning-based models are tightly coupled to specific input modalities or hardware (e.g., UWB) [22, 35], while optimization frameworks often suffer from high computational cost or require detailed geometric priors. Moreover, both approaches generally assume access to extensive measurement data, which is impractical in scenarios involving non-cooperative sources or inaccessible environments. In contrast, our work takes a fundamentally different approach by recasting the NLoS localization problem as a generative inference task, where the unknown radiation source location and signal distribution are sampled from a learned conditional distribution. By leveraging diffusion-based generative modeling, we learn the spatial prior distribution of electromagnetic signal propagation from environmental geometry and then perform posterior inference based on sparse RSS observations. This framework not only eliminates the need for pilot signals or pre-classified measurements but also offers theoretical guarantees on posterior recoverability under sparse sampling guided by diffraction-informed heuristics. In doing so, our method bridges the gap between data-driven modeling and physics-aware inference, offering a scalable and generalizable solution for high-precision NLoS localization in real-world, measurement-limited scenarios.
Figure 2: Illustration of the RadioDiff-Loc framework. The RSS information in the allowed areas and the full environment information are input to the DM as the condition to generate the full RM for NLoS localization.